Find all numbers for which the rational expression is undefined.
q^3-2q / q^2-4

Answer by rapaljer(4671)

A rational expression (fraction!) is undefined if the denominator is ZERO! In this case,
q^2-4 =0
(q-2)(q+2)=0
q=2 and q=-2

R^2